A trail mix comprised of 66% peanuts is being mixed with a trail mix variety that is 73% peanuts to create a 28 lb. bag of trail mix that is 68% peanuts. How many pounds of the 66% peanuts trail mix must be used?
Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39618)

Maybe this demonstration will help, although it is a different example, but of the same kind:

two-part mixture using one variable

Answer by stanbon(75887)

A trail mix comprised of 66% peanuts is being mixed with a trail mix variety that is 73% peanuts to create a 28 lb. bag of trail mix that is 68% peanuts. How many pounds of the 66% peanuts trail mix must be used?
------
peanuts + peanuts = peanuts
----
0.66x + 0.73(28-x) = 0.68*28
-------
66x + 73*28 - 73x = 68*28
-------
-7x = -5*28
----
x = 5*4 = 20 lb (amt. of 66% peanuts needed)
-----------
Cheers,
Stan H.
-----------